1. What would be the boiling point of liquid L at the top of a mountain where the

atmospheric pressure is lower than 1 atm?

a. 78.5 °C

b. less than 78.5 °C

c. greater than 78.5 °C

d. It depends on the amount of heat the liquid is exposed to.

2. The vapor pressure of ethyl alcohol at 60 °C is 47.02 kPa. What would be its vapor

pressure at 20 °C?

a. 47.02 kPa

b. greater than 47.02 kPa

c. less than 47.02 kPa

d. cannot be determined



1
Expert's answer
2021-03-31T08:01:28-0400

1).b). less than 78.5 °C


2)."\frac {P_1}{T_1} =\frac {P_2} {T_2}"


"P_2=(\frac {47.02×293}{333}) =41.372kPa"


c. less than 47.02 kPa